# Universality of Estimator for High-Dimensional Linear Models with Block Dependency

### v1

@@ Abstract.
We study the universality property of estimators for high-dimensional linear models, which exhibits a distribution of estimators is independent of whether covariates follows a Gaussian distribution. Recent high-dimensional statistics require covariates to strictly follow a Gaussian distribution to reveal precise properties of estimators. To relax the Gaussianity requirement, the existing literature has studied the conditions that allow estimators to achieve universality. In particular, independence of each element of the high-dimensional covariates plays an important role. In this study, we focus on high-dimensional linear models and covariates with block dependencies, where elements of covariates only within a block can be dependent, then show that estimators for the model maintain the universality. Specifically, we prove that a distribution of estimators with Gaussian covariates is approximated by an estimator with non-Gaussian covariates with same moments, with the setup of block dependence. To establish the result, we develop a generalized Lindeberg principle to handle block dependencies and derive new error bounds for correlated elements of covariates. We also apply our result of the universality to a distribution of robust estimators.
@@ 1. Introduction
We consider an estimator for a linear model by minimizing an empirical risk. Let $p\in\mathbb{N}$ be dimension and $\mathcal{Y}$ is a space of output variables. Let $n\in\mathbb{N}$ be a number of observations, and $(X_{1},Y_{1}),...,(X_{n},Y_{n})$ be a set of independently and identically distributed $\mathbb{R}^{p}\times\mathcal{Y}$-valued random observations generated from a statistical model.
@@
where $\theta_{0}\in\mathbb{R}^{p}$ is a unknown parameter and $\xi_{i}$ is an independent noise variable. With a convex loss function $\ell_{0}:\mathbb{R}\to\mathbb{R}$ and a convex regularizer $R_{0}:\mathbb{R}^{p}\to\mathbb{R}$, we consider an estimator $\widehat{\theta}_{\mathbf{X}}$ through the empirical risk minimization problem:
@@
In this study, we show the universality of the distribution of this estimator $\widehat{\theta}_{\mathbf{X}}$, that is, a distribution of the estimator does not depend on whether the covariate $X_{i}$ follow a Gaussian distribution. Specifically, we will show that the universality holds even in the presence of weak dependencies between elements of the random covariate $X_{i}$.
@@
The Gaussianity of covariates is a particularly important property in the analysis of high-dimensional statistics, which has been rapidly developing in recent years. Here, the term high-dimensional refers to the so-called proportional high-dimensional regime, where both the number of parameters $p$ in the statistical model and the amount of data $n$ used for estimation diverge while maintaining a certain ratio. In this setting, several advanced frameworks allow the precise determination of the risk and asymptotic distribution of estimators. These tools include not only the random matrix theory (Marchenko and Pastur, 1967; Dobriban and Wager, 2018; Hastie et al., 2022) but also frameworks such as the convex Gaussian minimax theorem (Stojnic, 2013; Thrampoulidis et al., 2015, 2018; Miolane and Montanari, 2021; Salehi et al., 2019; Montanari et al., 2019; Loureiro et al., 2021; Liang and Sur, 2022; Celentano et al., 2023; Koehler et al., 2021; Tsuda and Imaizumi, 2023), the approximate massage passing (Donoho et al., 2009; Bayati and Montanari, 2011a, b; Rangan, 2011; Javanmard and Montanari, 2013; Donoho and Montanari, 2016; Sur and Candès, 2019; Barbier et al., 2019; Mondelli and Venkataramanan, 2021; Zhao et al., 2022; Feng et al., 2022; Sawaya et al., 2023), the leave-one-out (El Karoui et al., 2013; Bean et al., 2013; El Karoui, 2018; Sur et al., 2019; Yadlowsky et al., 2021), and others (Mézard et al., 1987; Mezard and Montanari, 2009; Bellec, 2022; Sawaya et al., 2024), which have been applied to analyze the properties of statistical models across a wide range of domains. Although these developments are remarkable, many of these frameworks heavily rely on the strict Gaussianity of covariates, which poses a practical limitation.
@@
Relaxing the Gaussianity of covariates of these high-dimensional statistics is crucial to enhance the practicality of high-dimensional methods. The property that the behavior of statistics or estimators does not rely on Gaussianity is known as universality. Specifically, the universality means that properties of statistics under non-Gaussian covariates asymptotically matches that under Gaussian covariates with the same moments. Several studies (Korada and Montanari, 2011; Montanari and Nguyen, 2017; Hu and Lu, 2022; Goldt et al., 2022; Montanari and Saeed, 2022; Dudeja et al., 2022; Lahiry and Sur, 2024) have shown that, in a risk minimization problem, a value of the minimized risk using high-dimensional models achieves universality. The universality of approximate massage passing algorithms is also extensively studied using a rotationally invariant matrix (Fan, 2022; Dudeja et al., 2023; Wang et al., 2024). The universality has been studied not only in terms of risks but also in detailed properties of estimators required for statistical inference. Han and Shen (2023) shows that the estimators of high-dimensional parameters obtained by risk minimization achieve the universality. Han (2024) demonstrates that an estimator given by a first-order algorithm achieves the universality. Furthermore, the universality has been shown to hold in more specific high-dimensional models.
@@
A challenge in the universality of estimators lies in the requirement on properties of non-Gaussian covariates, particularly the independence between elements of covariates. In Han and Shen (2023) and Han (2024), it is stated that, for the universality of estimators, each element of high-dimensional covariates used for the estimator should be independent. This condition not only fails to hold in real high-dimensional covariates but also poses an obstacle when applying universality results to various statistical models. Related to this issue, Lahiry and Sur (2024) demonstrates that the universality of risks holds under the assumption that elements of $p$-dimensional covariates have weak dependence, specifically block dependence. However, this relaxation of the independence of elements for the universality of estimators remains an ongoing challenge.
@@
In this study, we show that the universality holds for estimators obtained by risk minimization, when elements of covariates exhibit dependencies. Specifically, we consider block dependency of elements of covariates, where each element of covariate vectors depends on at most $d$ other elements. We show that when the dependency coefficient $d$ has a certain polynomial order in $n$ or $p$, a distribution of estimators does not rely on the Gaussianity of covariates, in the limit where $n$ and $p$ diverge while maintaining a proportional relationship. This research follows the universality framework of estimators established by Han and Shen (2023), which allows the dependency considered in Lahiry and Sur (2024). As an application of the result, we demonstrate that we derive the universality of a distribution of robust estimators obtained by minimizing the absolute loss and the Huber loss.
@@
As technical aspects, we develop two proof techniques to demonstrate the universality results above. The first is a generalized Lindeberg principle tailored to address the block dependence. In our approach, we measure the discrepancy of two design matrices by focusing on a cell of dependent elements of covariates. We then show that this error is bounded by a cubic polynomial of the number of correlated elements under block dependence, which is critical to controlling the order of block dependence. The second technique involves deriving an upper bound on the number of dependent elements under block dependence. We generate a refined partition based on the original partition of elements, then we derive a tighter upper bound from the new partition. These two techniques allow for a precise evaluation of errors compared to the previous studies, thereby enabling the universality under block dependence with the larger order.
We mention the differences between our research and related studies. Lahiry and Sur (2024) is the closest to our work, which develops the notion of block dependence and derives the universality of risks under block dependence. While Lahiry and Sur (2024) utilizes Stein’s lemma to show the universality of risks, we developed a proof using the Lindeberg approach which enables us to show the universality of estimators. By this approach, we show the universality of the distribution of specific estimators, such as the robust estimator.
@@ 1.1. Notation
[…]
@@ 2.1. Setup: Estimator with Convex Risk Minimization
We consider an estimator with empirical loss minimization under general convex loss and convex regularization. Recall that the random $n$ observations $(X_{1},Y_{1}),...,(X_{n},Y_{n})$ follow a statistical model. Also, we assume that the $\mathbb{R}^{p}$-valued random vectors $X_{1},...,X_{n}$ follows $\mathbb{E}[X_{i}]=0$ and $\mathbb{E}[X_{i}X_{i}^{\top}]=\Sigma/n$ with a positive definite matrix $\Sigma\in\mathbb{R}^{p\times p}$. Also, recall that the estimator $\widehat{\theta}_{\mathbf{X}}$ defined in (2) with the convex loss function $\ell_{0}$ and the regularizer $f_{0}$. We further define design matrices $\mathbf{X}=(X_{1},...,X_{n})^{\top}\in\mathbb{R}^{n\times p}$ of the covariates.
@@ 2.2. Basic Assumption
This assumption restricts the high-dimensional setting we consider to the proportionally high-dimensional regime. Since $\tau$ ranges in an open interval $(0,1)$, we can also consider the setting as $p<n$.
@@
There exist non-negative real numbers $q_{\ell}$ ($\ell=0,1,2,3$), constants $\bar{\rho}\in(0,1]$, and ${L_{\ell_{i}}\geq 1:i\in[m]}$, along with two measurable functions $\mathcal{D}_{\ell}(\rho)$ and $\mathcal{M}_{\ell}(\rho):(0,\bar{\rho})\rightarrow\mathbb{R}_{\geq 0}$, where $\mathcal{M}_{\ell}(\rho)\leq 1\leq\mathcal{D}_{\ell}(\rho)$, that satisfy the following conditions:
@@
    The functions $\ell_{i}$ grow at most polynomially, meaning for $i\in[m]$,
@@
    There exist smooth approximations $\{\ell_{i;\rho}:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}\}_{\rho\in(0,\bar{\rho})}$ of $\ell_{i}$ such that (i) $\ell_{i;\rho}\in C^{3}(\mathbb{R})$, (ii) $\max_{i\in[m]}\|\ell_{i;\rho}-\ell_{i}\|_{\infty}\leq\mathcal{M}_{\ell}(\rho)$, and (iii) the derivatives of ${\ell_{i;\rho}}$ adhere to the following self-bounding property:
@@
Assumption 2 is exactly the same as an assumption on loss functions in Han and Shen (2023). $\mathcal{M}_{\ell}(\rho)$ and $\mathcal{D}_{\ell}(\rho)$ prescribe the smooth approximation of a loss function. In particular, when a loss function is in $C^{3}(\mathbb{R})$, $\mathcal{M}_{\ell}(\rho)$ is equal to 00. Assumption 2-(i) prevents loss functions from growing too fast as $|x|$ goes to infinity. In many cases, $L_{\ell_{i}}$ is characterized by a noise term of a statistical model. Assumption 2-(ii) restricts on the smoothness of loss functions. It allows loss functions to have a smooth approximation with respect to the supremum norm that increases at most polynomially as in Assumption 2-(ii) . We provide several common examples and their corresponding parameters $q$:
@@
Let $q^{*}$ equal to $2^{(\bar{\mathbf{q}}_{0}+4)/2}$ and assume there exists $X_{0}$ such that $X_{i}=X^{0}_{i}/\sqrt{n}$ and the following inequality hold:
$M:=\max$$\left\{\max_{1\leq j\leq p}\|X^{0}_{i,j}\|_{q^{*}},\lambda_{\max}(\Sigma)\right\}<\infty.$ (5)
In the universality studies, controlling the moments and the boundedness of the largest eigenvalue are both common assumptions (Montanari and Saeed, 2022; Han and Shen, 2023; Lahiry and Sur, 2024).
@@ 2.3. Dependence structure
In the following sections, we consider the situation where the observations $X_{1},...,X_{n}$ follows the distribution which satisfies the $d$-block dependence. Note that the observations are independently and identically distributed, hence the partition $\{B_{j}\}_{j=1}^{k}$ is independent of $i=1,...,n$.
@@
In several high-dimensional studies, block dependence and related concepts have been utilized. Han and Shen (2023) and Han (2024), which deal with the universality of estimators, assumes the independence of all elements, which can be considered as block dependence with $d=1$. The block dependence with $d\geq 2$ was introduced in Lahiry and Sur (2024), demonstrating universality under $d=o((n/\log n)^{1/5})$. In the field of high-dimensional central limit theorems, Chang et al. (2024) introduces $m$-dependence, a related notion of the block dependence.

### v2

# Universality of estimators for high-dimensional linear models with block dependency

@@ Abstract.
We study the universality property of estimators for high-dimensional linear models, which implies that the distribution of estimators is independent of whether the covariates follow a Gaussian distribution. Recent developments in high-dimensional statistics typically require covariates to strictly follow a Gaussian distribution to precisely characterize the properties of estimators. To relax this Gaussianity requirement, the existing literature has examined conditions under which estimators achieve universality. In particular, independence among the elements of the high-dimensional covariates has played a critical role. In this study, we focus on high-dimensional linear models with covariates exhibiting block dependence, where covariate elements can only be dependent within each block, and show that estimators for such models retain universality. Specifically, we prove that the distribution of estimators with Gaussian covariates can be approximated by the distribution of estimators with non-Gaussian covariates having the same moments under block dependence. To establish this result, we develop a generalized Lindeberg principle suitable for handling block dependencies and derive new error bounds for correlated covariate elements. We further demonstrate the universality result across several different estimators.
@@ 1. Introduction
We consider an estimator for a linear model obtained by minimizing an empirical risk. Let $p\in\mathbb{N}$ denote the dimension and let $\mathbb{R}$ denote a space of output variables. Let $n\in\mathbb{N}$ be the number of observations, and $(X_{1},Y_{1}),...,(X_{n},Y_{n})$ be a set of independently and identically distributed $\mathbb{R}^{p}\times\mathbb{R}$-valued random observations generated from a statistical model:
@@
where $\theta_{0}\in\mathbb{R}^{p}$ is an unknown parameter and $\xi_{i}$ are independent noise variables. Given a convex loss function $\ell_{0}:\mathbb{R}\to\mathbb{R}$ and a convex regularizer $R_{0}:\mathbb{R}^{p}\to\mathbb{R}$, we consider the estimator $\widehat{\theta}_{\mathbf{X}}$ via the empirical risk minimization problem:
@@
In this study, we show the universality of the distribution of this estimator $\widehat{\theta}_{\mathbf{X}}$, that is, the distribution of the estimator does not depend on whether the covariates $X_{i}$ follow a Gaussian distribution. Specifically, we will show that the universality holds even in the presence of weak dependencies among elements of the random covariate $X_{i}$.
@@
The Gaussianity of covariates is a particularly important property in the analysis of high-dimensional statistics, which has been rapidly developing in recent years. Here, the term high-dimensional refers to the so-called proportional high-dimensional regime, where both the number of parameters $p$ in the statistical model and the amount of data $n$ used for estimation diverge while maintaining a certain ratio. In this setting, several advanced frameworks allow the precise determination of the risk and asymptotic distribution of estimators. These tools include not only the random matrix theory (Marchenko and Pastur, 1967; Dobriban and Wager, 2018; Hastie et al., 2022) but also frameworks such as the convex Gaussian minimax (CGMT) theorem (Stojnic, 2013; Thrampoulidis et al., 2015, 2018; Miolane and Montanari, 2021; Salehi et al., 2019; Montanari et al., 2025; Loureiro et al., 2021; Liang and Sur, 2022; Celentano et al., 2023; Koehler et al., 2021; Tsuda and Imaizumi, 2024), the approximate message passing (Donoho et al., 2009; Bayati and Montanari, 2011a, b; Rangan, 2011; Javanmard and Montanari, 2013; Donoho and Montanari, 2016; Sur and Candès, 2019; Barbier et al., 2019; Mondelli and Venkataramanan, 2021; Zhao et al., 2022; Feng et al., 2022; Sawaya et al., 2023), the leave-one-out (El Karoui et al., 2013; Bean et al., 2013; El Karoui, 2018; Sur et al., 2019; Yadlowsky et al., 2021), and others (Mézard et al., 1987; Mezard and Montanari, 2009; Bellec, 2025; Sawaya et al., 2024), which have been applied to analyze the properties of statistical models across a wide range of domains. Although these developments are remarkable, many of these frameworks heavily rely on the strict Gaussianity of covariates, which poses a practical limitation.
@@
Relaxing the Gaussianity assumption on covariates in high-dimensional statistics is crucial for enhancing the practicality of high-dimensional methods. The property whereby the behavior of statistics or estimators does not depend on Gaussianity is known as universality. Specifically, universality implies that properties of statistics under non-Gaussian covariates asymptotically match those under Gaussian covariates sharing the same moments. Several studies (Korada and Montanari, 2011; Montanari and Nguyen, 2017; Hu and Lu, 2022; Goldt et al., 2022; Montanari and Saeed, 2022; Dudeja et al., 2024; Lahiry and Sur, 2024) have demonstrated that, in risk minimization problems, the minimized risk values using high-dimensional models exhibit universality. The universality of approximate message passing algorithms has also been extensively studied in the context of rotationally invariant matrices (Fan, 2022; Dudeja et al., 2023; Wang et al., 2024). Moreover, universality has been examined not only in terms of risks but also in terms of detailed properties of estimators required for statistical inference. Han and Shen (2023) shows that estimators for high-dimensional parameters obtained through risk minimization exhibit universality. Han (2025) demonstrates universality for estimators derived from first-order algorithms. Additionally, universality has been shown to hold within more specific high-dimensional models.
@@
A challenge in the universality of estimators lies in the conditions required for non-Gaussian covariates, particularly the independence among their elements. Han and Shen (2023) and Han (2025) state that, for estimators to exhibit universality, each element of the high-dimensional covariates must be independent. However, this condition often fails to hold for real high-dimensional covariates, posing an obstacle to applying universality results in various statistical models. To address this issue, Lahiry and Sur (2024) demonstrates that the universality of estimators and risks holds under weaker assumptions, specifically, block dependence among elements of $p$-dimensional covariates with a diagonal covariance matrix. Despite these advancements, relaxing independence for universal estimators remains a significant challenge.
@@
In this study, we show that universality holds for estimators obtained by risk minimization even when elements of covariates exhibit dependencies. Specifically, we consider block dependence, where each element of the covariate vectors depends on at most $d$ other elements. We demonstrate that when the dependency coefficient $d$ has a certain polynomial order in $n$ or $p$, the distribution of estimators does not depend on the Gaussianity of covariates in the limit where $n$ and $p$ diverge while maintaining a proportional relationship. Our research extends the universality framework established by Han and Shen (2023), incorporating the dependency structure considered in Lahiry and Sur (2024). As an application of our results, we establish universality for the distribution of robust estimators obtained by minimizing the absolute loss and the Huber loss, as well as of the lasso and ridge estimators.
@@
As technical aspects, we develop the following proof technique to demonstrate the universality results presented above. Specifically, we introduce a generalized Lindeberg principle tailored to handle block dependence. In our approach, we measure the discrepancy between two design matrices by focusing on cells of dependent covariate elements. We then show that this error is bounded by a cubic polynomial in the number of correlated elements under block dependence, which is crucial for controlling the order of block dependence. This technique enables a more precise evaluation of errors compared to previous studies, thus facilitating universality under higher-order block dependence.
We highlight the differences between our research and related studies. Lahiry and Sur (2024) is closest to our work, introducing the notion of block dependence and deriving universality results for estimators and risks under some dependence. A key difference of ours from Lahiry and Sur (2024) is the universality under general covariance structure. This contribution is based on the leverage of the new CGMT by Akhtiamov et al. (2025). Another difference is the proof technique: while Lahiry and Sur (2024) employs Stein’s lemma to establish the universality, we develop a proof based on the Lindeberg method.
@@ 1.1. Notation
For $x,x^{\prime}\in\mathbb{R}$, we define their maximum as $x\vee x^{\prime}:=\max\{x,x^{\prime}\}$. For $z\in\mathbb{N}$, $z]:=\{1,2,...,z\}$ is a set of natural numbers no more than $z$. For a vector $x=(x_{1},...,x_{d})^{\top}\in\mathbb{R}^{d}$, $x_{j}$ denotes the $j$-th element of $x$ for $j\in[d]$. For a random variable $X$ and $q\geq 1$, $\|X\|_{q}:=E[|X|^{q}]^{1/q}$ denotes its $q$-norm. For a positive semidefinite matrix $A$, $\|x\|_{A}^{2}:=\langle x,Ax\rangle$ denotes the Mahalanobis (semi-)norm. For $x\in\mathbb{R}^{d}$ and a function $h:\mathbb{R}^{d}\to\mathbb{R}$, $\partial_{j}h(x):=\partial h(x)/\partial{x_{j}}$ denotes a partial derivative in terms of $x_{j}$ for $j\in[d]$. For a matrix $A\in\mathbb{R}^{d_{1}\times d_{2}}$ and a function $h:\mathbb{R}^{d_{1}\times d_{2}}\to\mathbb{R}$, $\partial_{j_{1},j_{2}}h(A):=\partial h(A)/\partial{A_{j_{1},j_{2}}}$ denotes a partial derivative in terms of a $(j_{1},j_{2})$-element $A_{j_{1},j_{2}}$ of $A$. For a function $f:\Omega\to\mathbb{R}$ with $\Omega\subset\mathbb{R}^{d}$, $\|f\|_{L^{q}}:=(\int_{\Omega}|f(x)|^{q}dx)^{1/q}$ denotes the $L^{q}$-norm with $q\geq 1$. For a variable $x$, $C_{x}$ denotes a positive constant depending polynomially on $x$, whose value may change from line to line. In addition, $a\lesssim_{x}b$ denotes $a\leq C_{x}b$, and $a\lesssim b$ denotes $a\leq Cb$ with a universal constant $C>0$. $a\asymp b$ denotes both $a\lesssim b$ and $b\lesssim a$ holds. For a matrix $A\in\mathbb{R}^{d\times d}$, $\lambda_{\max}(A)$ denotes the largest eigenvalue of $A$. We define the Moreau envelope of a function $f:\mathbb{R}\to\mathbb{R}$ as $e_{f}(x;\lambda):=\min_{z}(2\lambda)^{-1}\|x-z\|_{2}^{2}+f(z)$ with $\lambda>0$.
@@ 2.1. Setup: estimator with convex risk minimization
We consider an estimator with empirical loss minimization under general convex loss and convex regularization. Recall that the random $n$ observations $(X_{1},Y_{1}),...,(X_{n},Y_{n})$ follow a statistical model. We assume that the $\mathbb{R}^{p}$-valued random vectors $X_{1},...,X_{n}$ follow $\mathbb{E}[X_{i}]=0$ and $\mathbb{E}[X_{i}X_{i}^{\top}]=\Sigma/n$ with a positive definite matrix $\Sigma\in\mathbb{R}^{p\times p}$. Also, recall that the estimator $\widehat{\theta}_{\mathbf{X}}$ is defined in (2) with the convex loss function $\ell_{0}$ and the regularizer $f_{0}$. We further define design matrices $\mathbf{X}=(X_{1},...,X_{n})^{\top}\in\mathbb{R}^{n\times p}$ of the covariates.
@@ 2.2. Basic assumptions
This assumption restricts the high-dimensional setting we consider to the proportionally high-dimensional regime.
@@
There exist non-negative real numbers $q_{\ell}$ ($\ell=0,1,2,3$), constants $\bar{\rho}\in(0,1]$, and $L_{\ell_{i}}\geq 1:i\in[n]$}, along with two measurable functions $\mathcal{D}_{\ell}(\rho)$ and $\mathcal{M}_{\ell}(\rho):(0,\bar{\rho})\rightarrow\mathbb{R}_{\geq 0}$, where $\mathcal{M}_{\ell}(\rho)\leq 1\leq\mathcal{D}_{\ell}(\rho)$, that satisfy the following conditions:
@@
    The functions $\ell_{i}$ grow at most polynomially, meaning for $i\in[n]$,
@@
    There exist smooth approximations $\{\ell_{i;\rho}:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}\}_{\rho\in(0,\bar{\rho})}$ of $\ell_{i}$ such that (a) $\ell_{i;\rho}\in C^{3}(\mathbb{R})$, (b) $\max_{i\in[m]}\|\ell_{i;\rho}-\ell_{i}\|_{\infty}\leq\mathcal{M}_{\ell}(\rho)$, and (c) the derivatives of $\ell_{i;\rho}$} adhere to the following self-bounding property:
@@
Assumption 2 is exactly the same as an assumption on loss functions in Han and Shen (2023). $\mathcal{M}_{\ell}(\rho)$ and $\mathcal{D}_{\ell}(\rho)$ prescribe the smooth approximation of a loss function. In particular, when a loss function is in $C^{3}(\mathbb{R})$, $\mathcal{M}_{\ell}(\rho)$ is equal to 0. Assumption 2-(i) prevents loss functions from growing too fast as $|x|$ goes to infinity. In many cases, $L_{\ell_{i}}$ is characterized by a noise term of a statistical model. Assumption 2-(ii) restricts the smoothness of loss functions. It allows loss functions to have a smooth approximation with respect to the supremum norm that increases at most polynomially as in Assumption 2-(ii) . We provide several common examples and their corresponding parameters $q$:
@@
Let $q^{*}$ equal to $2^{(\bar{\mathbf{q}}_{0}+4)/2}$ and assume there exists $X^{0}$ such that $X_{i}=X^{0}_{i}/\sqrt{n}$ and the following inequality holds:
$M:=\max$ $\left\{\max_{1\leq j\leq p}\|X^{0}_{i,j}\|_{q^{*}},\lambda_{\max}(\Sigma)\right\}<\infty,\quad{\text{uniformly in $p$}}.$ (5)
The boundedness of the maximum eigenvalue is trivially satisfied if the case with $\Sigma=I$, which has been exploited in the existing universality studies (Montanari and Saeed, 2022; Han and Shen, 2023). In our setup, since we consider a general covariance matrix $\Sigma$, we explicitly assume boundedness of the maximum eigenvalue. A similar assumption is introduced in Lahiry and Sur (2024).
@@ 2.3. Dependence structure
In the following sections, we consider the situation where the observations $X_{1},...,X_{n}$ follow the distribution which satisfies the $d$-block dependence. Note that the observations are independently and identically distributed, hence the partition $\{B_{j}\}_{j=1}^{k}$ is independent of i=1,,ni=1,...,n.
@@
In several high-dimensional studies, block dependence and related concepts have been utilized. Han and Shen (2023) and Han (2025), which deal with the universality of estimators, assume the independence of all elements, which can be considered as block dependence with $d=1$. The block dependence with $d\geq 2$ was introduced in Lahiry and Sur (2024), demonstrating universality under $d=o((n/\log n)^{1/5})$. In the field of high-dimensional central limit theorems, Chang et al. (2024) introduces $m$-dependence, a related notion of the block dependence.